OpenStudy (anonymous):

i need help with the file i add pl look at it help me I WILL AWARD MEDAL!!!!

OpenStudy (anonymous):

i need help with the file i add pl look at it help me I WILL AWARD MEDAL!!!!

OpenStudy (anonymous):

i need help with the file i add pl look at it help me I WILL AWARD MEDAL!!!!

OpenStudy (anonymous):

You want to find the equation for a line that passes through the two points: (1996,173) and (2002,215). First of all, remember what the equation of a line is: y = mx+b Where: m is the slope, and b is the y-intercept First, let's find what m is, the slope of the line... So what we need now are the two points you gave that the line passes through. Let's call the first point you gave, (1996,173), point #1, so the x and y numbers given will be called x1 and y1. Or, x1=1996 and y1=173. Also, let's call the second point you gave, (2002,215), point #2, so the x and y numbers here will be called x2 and y2. Or, x2=2002 and y2=215. Now, just plug the numbers into the formula for m above, like this: m= \[\frac{ 215-173 }{ 2002 -1996 } \] or... m= 42 / 6 or... m=7 So, we have the first piece to finding the equation of this line, and we can fill it into y=mx+b like this: y=7x+b Now, what about b, the y-intercept? To find b, think about what your (x,y) points mean: (1996,173). When x of the line is 1996, y of the line must be 173. (2002,215). When x of the line is 2002, y of the line must be 215. Because you said the line passes through each one of these two points, right? Now, look at our line's equation so far: y=7x+b. b is what we want, the 7 is already set and x and y are just two "free variables" sitting there. We can plug anything we want in for x and y here, but we want the equation for the line that specfically passes through the two points (1996,173) and (2002,215). So, why not plug in for x and y from one of our (x,y) points that we know the line passes through? This will allow us to solve for b for the particular line that passes through the two points you gave!. You can use either (x,y) point you want..the answer will be the same: (1996,173). y=mx+b or 173=7 × 1996+b, or solving for b: b=173-(7)(1996). b=-13799. (2002,215). y=mx+b or 215=7 × 2002+b, or solving for b: b=215-(7)(2002). b=-13799. See! In both cases we got the same value for b. And this completes our problem. The equation of the line that passes through the points (1996,173) and (2002,215) is y=7x-13799

OpenStudy (anonymous):

(b) use the line to estimate the population in 2003. the populationin 2003 is estimate to be _____ thousand nmates.

OpenStudy (anonymous):

Set x = to 2003 in the equation we found. y = 7 ( 2003) - 13799 y= 222

OpenStudy (anonymous):

try got it not it

OpenStudy (anonymous):

Do they want point form? (2003,222)?

OpenStudy (anonymous):

how did you add the file?

OpenStudy (anonymous):

attach fflie it by draw

OpenStudy (anonymous):

MichaelBoocher no im lost

OpenStudy (anonymous):

Screen. Maybe 222,000?

OpenStudy (anonymous):

look at got and see what you think

OpenStudy (anonymous):

2013 BRO!

OpenStudy (anonymous):

292 i got a red x

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3D+7x+-+13799+%2C+x+%3D+2013 I dont know what to say... its the answer.